OpenStudy (jilie):

plz help There are (7^13)^3 ⋅ 7^0 strawberries in a field. What is the total number of strawberries in the field? 1 7^16 7^17 7^39

OpenStudy (princeharryyy):

7^13^3 * 7^0 = 7^(13*3+0) = 7^39

OpenStudy (princeharryyy):